OpenStudy (anonymous):

How do you set up this problem? A pump can empty a small pond in 4 hours. Another more efficient pump can do the job in 3 hours. How long would it take to empty the pond if both pumps were used?

OpenStudy (radar):

First off, you know it will be less than 3 hours. Use product over sum. (4 X 3)/7 12/7 1 5/7 hour if you want to convert to minutes , go ahead.

OpenStudy (anonymous):

Thank you!

OpenStudy (radar):

There is another way to work it, where 1/4 + 1/3=1/combined pump time 3/12 + 4/12= 1/x 1/4 is amount pumped in 1 hour 1/3 is amount pumped in 1 hour combined (add) is the amount the combined pumps in 1 hour.

OpenStudy (radar):

Multiply thruy by 12 getting 3+4=12/x 7/1 = 12/x or 7x=12, x=12/7

OpenStudy (radar):

Same answer, product over sum is the way to go.

OpenStudy (radar):

If there were more than 2 pumps say 3, then do two using product over sum, and then use the result with the remaining one using product over sum.